Solve the ordinary differential equation
Y"+y''+y=0
1
Expert's answer
2020-04-28T13:31:27-0400

The question on this form is incorrect if the first term is y''' , the auxiliary equation does not has solution . So i will solve the equation on the form

"y''+y'+y=0"

Set up and solve the auxiliary equation

"\begin{aligned}m^2+m+1&=0\\
\end{aligned}"

Then

"m=\frac{-1\pm \sqrt{1-4}}{2}=\frac{-1}{2}\pm \frac{\sqrt{3}}{2}i"

Here "\alpha =-1/2,\ \ \ \beta= \sqrt{3}/2" , then the solution given by

"\begin{aligned}
y&=e^{\alpha x} [c_1\cos \beta x+c_2\sin \beta x]\\
&= e^{(-1/2)x}[c_1\cos (\sqrt{3}/2) x+c_2\sin(\sqrt{3}/2) x]
\end{aligned}"

---------------------------------------------------------------------------------------------------------------------------------------------

If the equation written true as

"y''+y''+y=0\ \ \to 2y''+y=0"

Set up and solve the auxiliary equation

"2m^2+1=0"

Then "m=\pm \frac{1}{\sqrt{2}}" , here "\alpha =0,\ \ \beta = 1/\sqrt{2}," then the solution given by

"\begin{aligned}
y&=e^{\alpha x} [c_1\cos \beta x+c_2\sin \beta x]\\
&= [c_1\cos (1/\sqrt{2}) x+c_2\sin(1/\sqrt{2}) x]
\end{aligned}"